A cat is walking at the speed of 5 m/s when he sees a mouse and begins to chase it.The cat speed up to a velocity of -4 m/s in 7 seconds. how far do the cat run to catch the mouse
1
Expert's answer
2020-10-23T11:51:04-0400

By definition, the acceleration is change in the velocity over the time, in which this change happens. Thus:


"a = \dfrac{v_2-v_1}{t} = \dfrac{-4m/s-5m/s}{7s} = -\dfrac{9}{7} m/s^2"

The distance travelled by the cat with this acceleration and in this time is:


"d = v_1t +\dfrac{at^2}{2} = 5m/s\cdot7s -\dfrac{9}{7}m/s^2\cdot\dfrac{(7s)^2}{2} =66.5m"

Answer. 66.5 m.